OpenStudy (anonymous):

Solve each system by substitution. x+7y=-13 -3x-3y+21

OpenStudy (richyw):

Substitution so solve for x or y in one equation. Usually either the easist one or the easiest one to "plug in" (next step).

OpenStudy (richyw):

wait, is the second one equal to 21?

OpenStudy (anonymous):

-3x-3y=21

OpenStudy (richyw):

the first one is easiest to do so we have \[x=-13-7y\]now just plug into the second equation \[-3(-13-7y)-3y=21\]\[39+21y-3y=21\]\[18y=-18\]\[y=-1\]now you can plug that value of y in to find x.

OpenStudy (richyw):

although doing it by "substitution" is way more work than you need to solve this basic system.